21x^2 -5x-6 how would i go about factoring this expression?
Answer by ankor@dixie-net.com(22740)

21x^2 -5x-6 how would i go about factoring this expression?
:
You just have to play around with the factors of 21 & 6
21: 3 & 7
6: 3 & 2
multiply factors
3 * 3 = 9
2 * 7 = 14
now give them the signs and position to equal -5
-14 + 9 = - 5
(3x - 2)(7x + 3), check is by FOILing